OpenStudy (anonymous):

write a fraction equivalent to 1/6 ?

OpenStudy (anonymous):

2/12

OpenStudy (minamoe):

2/12, 3/18, 4/24, 5/30, 6/36

OpenStudy (anonymous):

what about 1/2 ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

1/2 cant be = to 1/6

OpenStudy (anonymous):

no i mean an equivalent fraction to 1/2 ?

OpenStudy (anonymous):

50/100, 2/4

OpenStudy (anonymous):

what about 1/3 ? what equivalent

OpenStudy (anonymous):

2/6

OpenStudy (anonymous):

what about 2/5 ?

OpenStudy (anonymous):

4/10

OpenStudy (anonymous):

and 3/10 ..?

OpenStudy (dan815):

a fraction is equivalent if they can all be represented the same way in simplest terms

OpenStudy (dan815):

for example 2/12, is not the simplest as 2 divides the top and bottom (2/2 )/(12/2) = 1/6

OpenStudy (academicgurusinc):

The are many answers to this problem. One answer would be 2/12, another answer would be 4/24 and so on. You simply multiply denominator and numerator by 2. If you found this helpful, I encourage to subscribe to our youtube channel https://www.youtube.com/channel/UCYiI7SmkU4_vhdSzKBWsifg to stay current with all of our new videos. Regards, Academic Gurus Inc. Twitter (@Academic_Gurus) Facebook (AcademicGurusInc) Youtube (Academic Gurus Inc)